# Normal bases of small height in Galois number fields

### v1

@@ 1. Introduction and statement of results
Let $K/{\mathbb{Q}}$ be a quadratic extension. Then there exists a normal basis $\beta_{1},\beta_{2}\in K$ with
@@
We give a quick proof of this proposition in Section 2. Our first main result produces a general bound for any Galois extension $K/{\mathbb{Q}}$.
@@
$h(\beta_{i})\leq\frac{d^{4d}(d^{2}-d+2)^{4d-3}}{2}\binom{d-1}{[(d-1)/2]}^{2}|\Delta_{K}|^{(d-1)(4d-3)},$
@@
Let $K/{\mathbb{Q}}$ be a Galois extension of prime degree $d\geq 3$. There exists an integral normal basis $\beta_{1},\dots,\beta_{d}$ for $K$ over $\mathbb{Q}$} so that
@@ 2. Notation and heights
For any $\alpha\in{\mathcal{O}}_{K}$,
@@
for an absolute constant $c$. This implies that we must have $a\neq 0$, in which case $a\pm b\sqrt{D}$ is the desired normal basis. ∎
@@
We want to choose $\alpha\in{\mathcal{O}}_{K}$ so that $D(\alpha)\neq 0$. Let $\theta\in{\mathcal{O}}_{K}$ be a primitive element satisfying (2), then $1,\theta,\dots,\theta^{d-1}\in{\mathcal{O}}_{K}$ is a basis for $K$ over $\mathbb{Q}$} with
@@
$\leq$ $\frac{d^{4d}(d^{2}-d+2)^{4d-3}}{2}\binom{d-1}{[(d-1)/2]}^{2}|\Delta_{K}|^{(d-1)(4d-3)}.$
@@
Since we are assuming that $K/{\mathbb{Q}}$ is a Galois extension of prime degree $d\geq 3$, $K$ must be a totally real number field and so $d_{v}=1$ for every archimedean place $v$. Let $L_{K}=\Sigma_{K}({\mathcal{O}}_{K})$ be the full-rank lattice in ${\mathbb{R}}^{d}$, then each element ${\boldsymbol{\alpha}}=(\alpha_{1},\dots,\alpha_{d})\in L_{K}$ where the coordinates $\alpha_{1},\dots,\alpha_{d}$ are algebraic conjugates. Since $\alpha_{i}$’s are algebraic integers, we have $|\alpha_{i}|_{v}\leq 1$ for every $1\leq i\leq d$ and $v\nmid\infty$, hence, by product formula, for each ${\boldsymbol{\alpha}}\in L_{K}$,
@@
is a sublattice of rank d1d-1. Let us write ${\boldsymbol{\beta}}=(\beta_{1},\dots,\beta_{d})$ for the ${\boldsymbol{\alpha}}_{i}$ satisfying (14), then $\beta_{1},\dots,\beta_{d}\in{\mathcal{O}}_{K}$ is a normal basis for $K/{\mathbb{Q}}$ and (15) combined with (16) impliy that
@@
Acknowledgements: We wish to thank Professor Adebisi Agboola for a valuable suggestion that led us to consider an effective version of the normal basis theorem.

### v2

@@ 1. Introduction and statement of results
For all but at most finitely many quadratic extensions $K/{\mathbb{Q}}$, there exists a normal basis $\beta_{1},\beta_{2}\in K$ with
@@
We give a quick proof of this proposition in Section 2, showing that there can be at most finitely many exceptional quadratic fields $K/{\mathbb{Q}}$ with small discriminant for which this result does not hold. Our first main result produces a general bound for any Galois extension $K/{\mathbb{Q}}$.
@@
$h(\beta_{i})\leq\frac{d^{4d}(d^{2}-d+2)^{4d-3}}{2^{4d-3}}\binom{d-1}{[(d-1)/2]}^{2}|\Delta_{K}|^{(d-1)(4d-3)},$
@@
Let $K/{\mathbb{Q}}$ be a Galois extension of prime degree $d\geq 3$. There exists a normal basis $\beta_{1},\dots,\beta_{d}$ for $K$ over $\mathbb{Q}$} consisting of algebraic integers so that
@@ 2. Notation and heights
Let $K$ be a totally real number field. For any nonzero $\alpha\in{\mathcal{O}}_{K}$,
@@
We are only using the Minkowski embedding, ideal lattice construction and Lemma 2.5 in Section 4 where the $K:{\mathbb{Q}}]$ is an odd prime, which implies that $K$ is a totally real cyclic extension of $\mathbb{Q}$}. This is the reason why we are only introducing this notation in the totally real case, where it is simpler than in general.
@@
for an absolute constant $c$. This implies that we must have $a\neq 0$, in which case $a\pm b\sqrt{D}$ is the desired normal basis, unless $|\Delta_{K}|^{1/4}\leq c(2)/c$. Hence, there can be at most finitely many exceptions. ∎
@@
We want to choose $\alpha\in{\mathcal{O}}_{K}$ so that $D(\alpha)\neq 0$: if this is the case, then the conjugates of $g(\alpha)$ are $\mathbb{Q}$}-linearly independent, hence, form a normal basis for $K$ over $\mathbb{Q}$}. We proceed as follows. Let $\theta\in{\mathcal{O}}_{K}$ be a primitive element satisfying (2), then $1,\theta,\dots,\theta^{d-1}\in{\mathcal{O}}_{K}$ is a basis for $K$ over $\mathbb{Q}$} with
@@
$\leq$ $\frac{d^{4d}(d^{2}-d+2)^{4d-3}}{2^{4d-3}}\binom{d-1}{[(d-1)/2]}^{2}|\Delta_{K}|^{(d-1)(4d-3)}.$
@@
Since we are assuming that $K/{\mathbb{Q}}$ is a Galois extension of prime degree $d\geq 3$, then the order of the Galois group $G$ is an odd prime, so $G$ is cyclic; hence, $K/{\mathbb{Q}}$ is a cyclic extension. Further, $K$ must be a totally real number field and so $d_{v}=1$ for every archimedean place $v$. Therefore, the Minkowski embedding notation we introduced in Section 2 applies here. Let $L_{K}=\Sigma_{K}({\mathcal{O}}_{K})$ be the full-rank lattice in ${\mathbb{R}}^{d}$, then each element ${\boldsymbol{\alpha}}=(\alpha_{1},\dots,\alpha_{d})\in L_{K}$ where the coordinates $\alpha_{1},\dots,\alpha_{d}$ are algebraic conjugates. Since $\alpha_{i}$’s are algebraic integers, we have $|\alpha_{i}|_{v}\leq 1$ for every $1\leq i\leq d$ and $v\nmid\infty$, hence, by product formula, for each ${\boldsymbol{\alpha}}\in L_{K}$,
@@
is a sublattice of rank d1d-1. Let us write ${\boldsymbol{\beta}}=(\beta_{1},\dots,\beta_{d})$ for the ${\boldsymbol{\alpha}}_{i}$ satisfying (14), then $\beta_{1},\dots,\beta_{d}\in{\mathcal{O}}_{K}$ is a normal basis for $K/{\mathbb{Q}}$ and (15) combined with (16) imply that
@@
Acknowledgements: We wish to thank Professor Adebisi Agboola for a valuable suggestion that led us to consider an effective version of the normal basis theorem. We also thank the referee for a thorough review of our paper and helpful suggestions.
@@
Conflict of interest statement: The authors declare no conflict of interest.